# Parallel computations for Metropolis Markov chains with Picard maps

### v1

@@ 3.1 Assumption on the target
We study convergence of the Picard map under the following assumption on $\pi$.
Assumption 2 (Smooth and Hessian Lipschitz target).
The target distribution $\pi$ has a Lebesgue density in $\mathcal{X}=\mathbb{R}^{d}$ proportional to $\exp(-V)$ where $V$ is a twice-differentiable, $L$-smooth and $(\gamma L^{3/2}d^{1/2})$-strong Hessian-Lipschitz for some $\gamma\geq 0$, meaning that the Hessian of $V$ satisfies
@@
for some $M$ with $\mathrm{Tr}(M)\leq\gamma L^{3/2}d^{1/2}$, where $A\preccurlyeq B$ if $A-B$ is positive semi-definite for two symmetric matrices $A,B$.
The strong Hessian-Lipschitz condition implies the more classical Hessian-Lipschitz condition, $\|\nabla^{2}V(y)-\nabla^{2}V(x)\|\leq m^{\star}\|y-x\|$, where $m^{\star}=\lambda_{max}(M)$, while it is weaker than assuming a bounded Frobenius norm of the third derivative, as previously used for analysis of MCMC algorithms (cf. Chen and Gatmiry, 2023). Note also that the analysis of the convergence of $\Phi$ does not require strong convexity of $V$.
@@ 3.2 Metropolis-Hastings kernels
In this work, we focus on two popular zeroth-order Metropolis-Hasting Markov chains: Random Walk Metropolis (RWM) (Metropolis et al., 1953) and Metropolis within Gibbs (MwG) (Geman and Geman, 1984). Both satisfy Assumption 1 and can be simulated with the Online Picard algorithm. We start with the RWM case and defer the MwG one to Section 4.
Assumption 3 (RWM).
$X_{0},X_{1},\dots$ is a RWM Markov chain on $\mathcal{X}=\mathbb{R}^{d}$ with step-size $\xi=h/(\sqrt{Ld})$ for some $h>0$, i.e. it follows (1) with $W=(Z,U)$, $U\sim\mathrm{Unif}([0,1])$ uniformly distributed, $Z\sim\mathcal{N}(0,h^{2}I_{d}/(Ld))$ independently of $U$,
@@ 3.3 Main results
Theorem 1 (Probability of an incorrect guess).
Under Assumptions 2-3, for all $x_{0}\in\mathcal{X}$, $w_{0}\in\mathcal{W}$, we have
@@
Theorem 1 essentially states that, after a logarithmic number of Picard recursions, the probability that the Picard map correctly predicts the $i$th step is $\mathcal{O}(i/d)$.
Remark 1.
By Theorem 1, the probability of correctly guessing individual increments can be controlled up to $i=\mathcal{O}(d)$. Assuming (9) to be tight, this also suggests that increasing the number of parallel processors beyond $K=\mathcal{O}(d)$ provides little gains in terms of speeding-up convergence of the Picard recursion, as numerically observed in Section 6.
Remark 2.
The upper bound of Theorem 1 becomes vacuous for $i>d$. This is consistent with the fact that the mixing time of RWM under Assumptions 2 is $\mathcal{O}(d)$. Indeed, if $X_{0}$ and $X_{i}$ are approximately independent, then the prediction of $B(X_{i},U_{i},Z_{i})$ with $B(X_{0},U_{i},Z_{i})$ is uninformative.
By the union bound, Theorem 1 implies that, with high probability, the Online Picard algorithm makes $\mathcal{O}(\sqrt{d})$ consecutive correct guesses after $\mathcal{O}(\log(d))$ calls of the Picard map when using $K=\mathcal{O}(\sqrt{d})$ processors, as shown below.
Corollary 1.
Under Assumptions 2-3, for all $x_{0}\in\mathcal{X},\,w_{0}\in\mathcal{W}$, $\epsilon>0$, $K\leq\sqrt{\epsilon^{2}d/(2c_{0})}$, $c_{0}$ as in Theorem 1, $j=\lceil\log(d)\rceil$, $h\geq 1$, we have that
@@
Notice the validity of the bound in Corollary 1 for any starting value $X_{0}$ and innovation $W_{0}$. This allows to prove the next main theorem, where we bound with high probability the number of parallel iterations required by the Online Picard algorithm to simulate a given number of steps of RWM with $K=\mathcal{O}(\sqrt{d})$ parallel processors.
Theorem 2.
Suppose Assumptions 2-3 holds with $h\geq 1$. Then, for all $\delta>0$, $n\in\mathbb{N}$, $K\leq\sqrt{d/(8c_{0})}$, $c_{0}$ as in Theorem 1, we have that $\mathbb{P}(L^{(j)}>n)\geq 1-\delta$ with
$j=\lceil\log(d)\max\{4n/K,2\log(1/\delta)\}\rceil=\mathcal{O}\left(\frac{n}{K}\right)\,.$
By Theorem 2, the Online Picard algorithm achieves the optimal speedup relative to the sequential algorithm for $K$ up to $\mathcal{O}(\sqrt{d})$, since it simulates $n$ steps of the sequential algorithm with $\mathcal{O}(n/K)$ parallel iterations and $K$ parallel processors. Notice that the convergence of the Picard map does not deteriorate with the condition number $L/m$ of the target.
Combining Theorem 2 with available explicit bounds on the mixing times of RWM (Andrieu et al., 2024) gives the following complexity result.
Corollary 2 (Parallel round complexity of Online Picard algorithm).
Suppose Assumptions 2-3 holds with $mI_{d}\preccurlyeq\nabla^{2}V$, $h=1$ and $X_{0}\sim\mathcal{N}(x^{\star},L^{-1}I_{d})$, where $x^{\star}$ is the minimizer of $V$. Then, for any $\epsilon>0$, the Online Picard algorithm with $K=\lfloor\sqrt{d/(8c_{0})}\rfloor$, $c_{0}$ as in Theorem 1, outputs a random variable $X$, with $\|\mathcal{L}(X)-\pi\|_{\mathrm{TV}}\leq\epsilon$ after
$J=\mathcal{O}\left(\frac{L}{m}\sqrt{d}\,\mathrm{polylog}(\epsilon^{-1})\right)$
@@
Corollary 2 implies that the Online Picard algorithm with $K=\mathcal{O}(\sqrt{d})$ achieves a Parallel round complexity of $\mathcal{O}(\sqrt{d})$.
@@ 3.4 Contraction of the Picard map: comparison between first- and zeroth-order case
for a tuning parameter $\xi>0$. Under Assumption 2, for all $x,y\in\mathcal{X}^{K+1},$ with $x_{0}=y_{0}$, all $w\in\mathcal{W}^{K}$, and for all $i\leq K$, we have
$\|\Phi_{i}^{\mathrm{ULA}}(x,w)-\Phi_{i}^{\mathrm{ULA}}(y,w))\|$ $=\frac{\xi^{2}}{2}\|\sum_{\ell=0}^{i-1}(\nabla V(x_{\ell})-\nabla V(y_{\ell}))\|\leq\frac{\xi^{2}L}{2}\sum_{\ell=0}^{i-1}\|x_{\ell}-y_{\ell}\|,$
while Lemmas 5-6 in the supplement material imply that, under Assumptions 2-3,
@@
where $W\sim\nu^{\otimes K}$, with marginals as in (8), $c_{1}=(\sqrt{\frac{2}{\pi}}+\frac{\gamma\xi\sqrt{Ld}}{2})$ and $\delta(d)$ as in Theorem 1. Note that the contraction of the Picard map occurs pointwise (for each innovation $w$) for ULA and in expectation for RWM. More crucially, the RWM contraction has a slower rate, mainly due to the the square root function in (10).
3.5 Behavior in the tails
The convergence bounds shown in Section 3.3 apply to any initialization $X_{0}$. Here we show that the convergence is faster when the chain is in the tails of the distribution.
Proposition 2.
Suppose Assumptions 2-3 hold, $\ell I_{d}\preccurlyeq\nabla^{2}V$ and $X^{(0)}_{\ell}=x_{0}$ for all $\ell\geq 0$. Then, for all $j\geq 0$ and $K>0$, we have
$\lim_{|x_{0}|\to\infty}\mathbb{P}\left(L^{(j)}=jK\right)=1.$
Proposition 2 implies that, as the initial point goes further in tails, the convergence of the Picard map becomes faster (up to converging in one iteration). This is partly related to classical scaling limits obtained in the transient phase for RWM (Christensen et al., 2005), which show that that RWM in the tails has a deterministic behavior. In our context, Proposition 2 suggests that the Picard map converges faster in the transient phase, which we do observe in the numerical simulations of Section 6.3.
4 Metropolis within Gibbs
We now extend our convergence results to Metropolis within Gibbs (MwG) Markov chains. Our main motivation to consider MwG is that it displays better empirical performances in Section 6.
@@ 4.1 Metropolis within Gibbs kernels
For a given orthonormal basis $\{o_{0},o_{1},\dots,o_{d-1}\}$ spanning $\mathbb{R}^{d}$ (e.g. the standard basis), the (deterministic-scan) Metropolis within Gibbs (MwG) chain on $\mathbb{R}^{d}$ (Geman and Geman, 1984) performs iteratively a one-dimensional Metropolis-Hasting update in the subspace spanned by $o_{i}$, for $i=0,\dots,d-1$. The corresponding chain $(X_{i})_{i=0,1,\dots}$ can be treated as a time-inhomogeneous Markov chain, where every block of $d$ coordinates of the Markov chain is iteratively defined as
[…]
with innovations $W_{i}=(U_{i},\,Z_{i})$, where $Z_{0},\dots,Z_{d-1}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\nu$ are one dimensional symmetric random variables, $U_{0},\dots,U_{d-1}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathrm{Unif}([0,1])$, independently of $Z_{0},\dots,Z_{d-1}$ and
[…]
@@
4.2 Analysis
If we denote by $\mathrm{Haar}(d)$ the Haar measure on the orthogonal group in dimension $d$, our theoretical results hold with the following assumption.
Assumption 4.
$X_{0},X_{1},\dots$ is a MwG chain on $\mathcal{X}=\mathbb{R}^{d}$ with $(o_{0},o_{1},\dots,o_{d-1})\sim\mathrm{Haar}(d)$ and $Z=h/\sqrt{Ld}(2P-1)S$, $P\sim\mathrm{Ber}(1/2)$ and $S\sim\chi(d)$, where $\mathrm{Ber}(p)$ and $\chi(d)$ are respectively the Bernoulli distribution with parameter $p$ and the Chi distribution with $d$ degrees of freedom.
Theorem 3.
Under Assumptions 2 and 4,
  • the statement of Theorem 1 holds with $c_{0}$ and $\delta(d)$ replaced by $c_{0}=2h^{4}(\sqrt{\frac{2}{\pi}}+\frac{h\gamma}{2})^{2}$, $\delta(d)=11\exp(-d/10)$;
  • the statement of Theorem 2 holds with $h>5$, $c_{0}$ as above;
  • the statement of Proposition 4 below holds with $c_{0}$ as above and $d\geq-10\log(\epsilon r/33)$.
Theorem 3 shows that the analysis of the Picard map derived in Section 3.3 can be extended to MwG. However, for MwG, it is not possible to derive the parallel round complexity (as in Corollary 2) since, to the best of our knowledge, tight complexity bounds for MwG under a feasible start remain unknown, with the only available result known to us being a bound on the spectral gap provided in Ascolani et al. (2024, Corollary 7.4).
@@ 4.2 Analysis
Proposition 3 (MwG with isotropic Gaussian targets).
Suppose Assumption 4 holds and $V(x)=\|x-\mu\|^{2}/(2\sigma^{2})$, $\mu\in\mathbb{R}^{d},\sigma^{2}>0$. For every $K\leq d$ and $X_{0}\in\mathcal{X}$, we have $L^{(j)}=jK$.
Note that the result in Proposition 3 is deterministic, i.e. it holds for every sequence $w\in\mathcal{W}^{\infty}$.

### v2

@@ 3.1 Assumption on the target
We study convergence of the Picard map $\Phi$ under the following smoothness assumption on $\pi$. Note that the analysis of the convergence of $\Phi$ does not require strong convexity of $V$.
{assumption}
[Smooth and Hessian Lipschitz target] The target distribution $\pi$ has a Lebesgue density in $\mathcal{X}=\mathbb{R}^{d}$ proportional to $\exp(-V)$ where $V$ is a twice-differentiable, $L$-smooth and $(\gamma L^{3/2}d^{1/2})$-strongly Hessian-Lipschitz for some $\gamma\geq 0$, meaning that the Hessian of $V$ satisfies
@@
for some matrix $M$ with $\mathrm{Tr}(M)\leq\gamma L^{3/2}d^{1/2}$, where $A\preccurlyeq B$ if ABA-B is positive semi-definite for two symmetric matrices $A,B$. The strongly Hessian-Lipschitz condition implies the more classical Hessian-Lipschitz condition, $\|\nabla^{2}V(y)-\nabla^{2}V(x)\|\leq m^{\star}\|y-x\|$, where $m^{\star}=\lambda_{max}(M)$, while it is weaker than assuming a bounded Frobenius norm of the third derivative, as previously used for analysis of MCMC algorithms (cf. Chen & Gatmiry, 2023). This assumption holds, for example, for standard Bayesian logistic regression models with Gaussian priors, see Section A in the supplement for details. In general, we acknowledge that Assumption 3.1 can be restrictive and, while needed for the theoretical analysis, we do not expect it to be necessary for the proposed Picard algorithm to perform well in practice. For example, the numerical simulations in Section 6 suggest that our theoretical results can be predictive of practical performances beyond the assumptions required to derive them (see e.g. the examples on Poisson regression models therein).
@@ 3.2 Metropolis-Hastings kernels
In this work, we focus on two popular zeroth-order Metropolis-Hasting Markov chains: Random Walk Metropolis (RWM) (Metropolis et al., 1953) and Metropolis within Gibbs (MwG) (Geman & Geman, 1984). Both MCMC have piecewise constant Picard maps and can be simulated with the Online Picard algorithm. We start with the RWM case and defer the MwG case to Section 4.
{assumption}
[RWM] $X_{0},X_{1},\dots$ is a RWM Markov chain on $\mathcal{X}=\mathbb{R}^{d}$ with stepsize $\xi=h/(\sqrt{Ld})$ for some $h>0$, i.e. it follows (1) with $W=(Z,U)$, $U\sim\mathrm{Unif}([0,1])$ uniformly distributed, $Z\sim\mathcal{N}(0,h^{2}I_{d}/(Ld))$ independently of $U$,
@@ 3.3 Main results
Theorem 3.1 (Probability of an incorrect guess).
Under Assumptions 3.1-3.2, for all $x_{0}\in\mathcal{X}$, $w_{0}\in\mathcal{W}$, we have
@@
Theorem 3.1 essentially states that, after a logarithmic number of Picard recursions, the probability that the Picard map incorrectly predicts the $i$th step is $\mathcal{O}(i/d)$.
Remark 3.2.
By Theorem 3.1, the probability of correctly guessing individual increments can be controlled up to $i=\mathcal{O}(d)$. Assuming (6) to be tight, this also suggests that increasing the number of parallel processors beyond $K=\mathcal{O}(d)$ provides little gains in terms of speeding up convergence of the Picard recursion, as numerically observed in Section 6.
Remark 3.3.
The upper bound of Theorem 3.1 becomes vacuous for $i>d$. This is consistent with the fact that the mixing time of RWM under Assumptions 3.1 is $\mathcal{O}(d)$. Indeed, if $X_{0}$ and $X_{i}$ are approximately independent, then the prediction of $B(X_{i},U_{i},Z_{i})$ with $B(X_{0},U_{i},Z_{i})$ is uninformative.
By the union bound, Theorem 3.1 implies that, with high probability, the Online Picard algorithm makes $\mathcal{O}(\sqrt{d})$ consecutive correct guesses after $\mathcal{O}(\log(d))$ calls of the Picard map when using $K=\mathcal{O}(\sqrt{d})$ processors, as shown below.
Corollary 3.4.
Under Assumptions 3.1-3.2, for all $x_{0}\in\mathcal{X},\,w_{0}\in\mathcal{W}$, $\epsilon>0$, $K\leq\sqrt{\epsilon^{2}d/(2c_{0})}$, $c_{0}$ as in Theorem 3.1, $j=\lceil\log(d)\rceil$, $h\geq 1$, we have that
@@
Notice the validity of the bound in Corollary 3.4 for any starting value $X_{0}$ and innovation $W_{0}$. This allows to prove the next main theorem, where we bound with high probability the number of parallel iterations required by the Online Picard algorithm to simulate a given number of steps of RWM with $K$ up to $\mathcal{O}(\sqrt{d})$ parallel processors.
Theorem 3.5.
Suppose Assumptions 3.1-3.2 hold with $h\geq 1$. Then, for all $\epsilon>0$, $N\in\mathbb{N}$, $K\leq\sqrt{d/(8c_{0})}$, $c_{0}$ as in Theorem 3.1, we have that $\mathbb{P}(L^{(j)}>N)\geq 1-\epsilon$ with
$j=\left\lceil\log(d)\max\{4N/K,2\log(1/\epsilon)\}\right\rceil=\mathcal{O}\left(\frac{N}{K}\right)\,.$
By Theorem 3.5, the Online Picard algorithm achieves the optimal speedup relative to the sequential algorithm for $K$ up to $\mathcal{O}(\sqrt{d})$, since it simulates $N$ steps of the sequential algorithm with $\mathcal{O}(N/K)$ parallel iterations and $K$ parallel processors. Notice that the convergence of the Picard map does not deteriorate with the condition number $L/m$ of the target.
Combining Theorem 3.5 with available explicit bounds on the mixing times of RWM (Andrieu et al., 2024) gives the following complexity result.
Corollary 3.6 (Parallel round complexity of Online Picard algorithm).
Suppose Assumptions 3.1-3.2 hold with $mI_{d}\preccurlyeq\nabla^{2}V$, $h=1$ and $X_{0}\sim\mathcal{N}(x^{\star},L^{-1}I_{d})$, where $x^{\star}$ is the minimizer of $V$. Then, for any $\epsilon>0$, the Online Picard algorithm with $K\leq\lfloor\sqrt{d/(8c_{0})}\rfloor$, $c_{0}$ as in Theorem 3.1, outputs a random variable $X$, with $\|\mathcal{L}(X)-\pi\|_{\mathrm{TV}}\leq\epsilon$ after
$J=\mathcal{O}\left(\frac{L}{m}\frac{d}{K}\,\mathrm{polylog}(\epsilon^{-1})\right)$
@@
Corollary 3.6 implies that the Online Picard algorithm with $K=\mathcal{O}(\sqrt{d})$ achieves a speedup of $\mathcal{O}(\sqrt{d})$ compared to the sequential algorithm.
The convergence bounds in Theorem 3.5 apply to any initialization $X_{0}$. With the next proposition, we show that the convergence is faster when the chain is in the tails of the distribution.
Proposition 3.7.
Suppose Assumptions 3.1-3.2 hold, $mI_{d}\preccurlyeq\nabla^{2}V$ and $X^{(0)}_{\ell}=x_{0}$ for all $\ell\geq 0$. Then, for all $j\geq 0$ and $K>0$, we have
$\lim_{\|x_{0}\|\to\infty}\mathbb{P}\left(L^{(j)}=jK\right)=1.$
Proposition 3.7 implies that, as the initial point goes further into the tails, the convergence of the Picard map becomes faster (up to converging in one iteration). This is partly related to classical scaling limits obtained in the transient phase for RWM (Christensen et al., 2005), which show that RWM in the tails has a deterministic behavior. In our context, Proposition 3.7 suggests that the Picard map converges faster in the transient phase, which we do observe in the numerical simulations of Section 6.6.
@@ 3.4 Contraction of the Picard map: comparison between first- and zeroth-order case
for a tuning parameter $\xi>0$. Under Assumption 3.1, for all $x,y\in\mathcal{X}^{K+1},$ with $x_{0}=y_{0}$, all $w\in\mathcal{W}^{K}$, and for all $i\leq K$, we have
$\|\Phi_{i}^{\mathrm{ULA}}(x,w)-\Phi_{i}^{\mathrm{ULA}}(y,w)\|$ $=\frac{\xi^{2}}{2}\|\sum_{\ell=1}^{i-1}(\nabla V(x_{\ell})-\nabla V(y_{\ell}))\|\leq\frac{\xi^{2}L}{2}\sum_{\ell=1}^{i-1}\|x_{\ell}-y_{\ell}\|,$
while Lemmas B.9-B.11 in the supplementary material imply that, under Assumptions 3.1-3.2,
@@
where $W\sim\nu^{\otimes K}$, with marginals as in (5), $c_{1}=(\sqrt{\frac{2}{\pi}}+\frac{\gamma\xi\sqrt{Ld}}{2})$ and $\delta(d)$ as in Theorem 3.1. Note that the contraction of the Picard map occurs pointwise (for each innovation $w$) for ULA and in expectation for RWM. More crucially, the RWM contraction has a slower rate, mainly due to the square root function in (7).
4 Extension to Metropolis within Gibbs
@@ 4.1 Metropolis within Gibbs kernels
We now extend our convergence results to Metropolis within Gibbs (MwG) Markov chains. Our main motivation for considering MwG is that it displays better empirical performance in our numerical simulations (see Appendix E in the supplementary material).
For a given orthonormal basis $\{o_{0},o_{1},\dots,o_{d-1}\}$ spanning $\mathbb{R}^{d}$ (e.g. the standard basis), the (deterministic-scan) Metropolis within Gibbs (MwG) chain on $\mathbb{R}^{d}$ (Geman & Geman, 1984) performs iteratively a one-dimensional Metropolis-Hastings update in the subspace spanned by $o_{i}$, for $i=0,\dots,d-1$. The corresponding chain $(X_{i})_{i=0,1,\dots}$ can be treated as a time-inhomogeneous Markov chain, where every block of $d$ coordinates of the Markov chain is iteratively defined as
$X_{i+1}$ $=X_{i}+f_{i}(X_{i},W_{i}),$ $\text{ for }i=0,1,\dots,d-1,$
with innovations $W_{i}=(U_{i},\,Z_{i})$, where $Z_{0},\dots,Z_{d-1}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\nu$ are one-dimensional symmetric random variables, $U_{0},\dots,U_{d-1}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\mathrm{Unif}([0,1])$, independently of $Z_{0},\dots,Z_{d-1}$ and
$f_{i}(x,(U,Z))$ $=B(x,U,o_{i}Z)o_{i}Z\,$, $\text{ for }i=0,1,\dots,d-1,$ (8)
@@
4.2 Results for Metropolis within Gibbs
If we denote by $\mathrm{Haar}(d)$ the Haar measure on the orthogonal group in dimension $d$, our theoretical results hold with the following assumption. {assumption} $X_{0},X_{1},\dots$ is a MwG chain on $\mathcal{X}=\mathbb{R}^{d}$ with $(o_{0},o_{1},\dots,o_{d-1})\sim\mathrm{Haar}(d)$ and $Z=h/\sqrt{Ld}(2P-1)S$, $P\sim\mathrm{Ber}(1/2)$ and $S\sim\chi(d)$, where $\mathrm{Ber}(p)$ and $\chi(d)$ are respectively the Bernoulli distribution with parameter $p$ and the Chi distribution with $d$ degrees of freedom.
Theorem 4.1.
Under Assumptions 3.1 and 4.2,
  • the statement of Theorem 3.1 holds with $c_{0}$ and $\delta(d)$ replaced by $c_{0}=2h^{4}(\sqrt{\frac{2}{\pi}}+\frac{h\gamma}{2})^{2}$, $\delta(d)=11\exp(-d/10)$;
  • the statement of Theorem 3.5 holds with $h>5$, $c_{0}$ as above;
  • the statement of Proposition 5.1 below holds with $c_{0}$ as above and $d\geq-10\log(\epsilon r/33)$, where $r\geq 0$ is the tolerance error of the Approximate Picard defined in Section 5.
Theorem 4.1 shows that the analysis of the Picard map derived in Section 3.3 can be extended to MwG. However, for MwG, it is not possible to derive the parallel round complexity (as in Corollary 3.6) since, to the best of our knowledge, tight complexity bounds for MwG under a feasible start remain unknown, with the only available result known to us being a bound on the spectral gap provided in Ascolani et al. (2024, Corollary 7.4).
@@ 4.2 Results for Metropolis within Gibbs
Proposition 4.2 (MwG with isotropic Gaussian targets).
Suppose Assumption 4.2 holds and $V(x)=\|x-\mu\|^{2}/(2\sigma^{2})$, $\mu\in\mathbb{R}^{d},\sigma^{2}>0$. For every $K\leq d$ and $X_{0}\in\mathcal{X}$, we have $L^{(j)}=jK$.
Note that the result in Proposition 4.2 is deterministic, i.e. it holds for every sequence $w\in\mathcal{W}^{\infty}$.